Required information

[The following information applies to the questions displayed below.]

Widmer Watercraft’s predetermined overhead rate for the year 2017 is 200% of direct labor. Information on the company’s production activities during May 2017 follows.

  1. Purchased raw materials on credit, $240,000.
  2. Materials requisitions record use of the following materials for the month.
Job 136 $ 50,000
Job 137 32,500
Job 138 19,600
Job 139 22,600
Job 140 7,000
Total direct materials

131,700

Indirect materials 20,000
Total materials used $ 151,700
  1. Paid $15,500 cash to a computer consultant to reprogram factory equipment.
  2. Time tickets record use of the following labor for the month. These wages were paid in cash.
Job 136 $ 12,200
Job 137 10,600
Job 138 37,700
Job 139 39,000
Job 140 3,600
Total direct labor 103,100
Indirect labor 24,000
Total $ 127,100
  1. Applied overhead to Jobs 136, 138, and 139.
  2. Transferred Jobs 136, 138, and 139 to Finished Goods.
  3. Sold Jobs 136 and 138 on credit at a total price of $545,000.
  4. The company incurred the following overhead costs during the month (credit Prepaid Insurance for expired factory insurance).
Depreciation of factory building $ 68,500
Depreciation of factory equipment 37,500
Expired factory insurance 11,000
Accrued property taxes payable 37,000
  1. Applied overhead at month-end to the Work in Process Inventory account (Jobs 137 and 140) using the predetermined overhead rate of 200% of direct labor cost.

2. Prepare journal entries to record the events and transactions a through i.
Answer #1
No. Transaction. Account titles and explanation Debit Credit
1. a. Raw materials inventory $240000
Accounts payable $240000
(To record raw materials purchased on credit)
2. b. Work in process inventory $131700
Factory overhead $20000
Raw materials inventory $151700
(To record
3. c. Factory equipment $15500
Cash $15500
(To record factory equipment purchased)
4. d. Work in process inventory $103100
Factory overhead $24000
Cash $127100
(To record wages paid in cash)
5. e. Work in process inventory ($12200+37700+39000)*200% $177800
Factory overhead $177800
(To record applied overhead to Jobs 136, 138, and 139)
6. f. Finished goods inventory $358900
Work in process inventory $358900
(To record jobs transferred to finished goods)
7. g. Accounts receivable $545000
Sales revenue $545000
(To record jobs sold on credit)
Cost of goods sold $219300
Finished goods inventory $219300
(To record cost of goods sold)
8. h. Factory overhead $154000
Accumulated depreciation- Factory building $68500
Accumulated depreciation- Factory equipment $37500
Prepaid insurance $11000
Property tax payable $37000
(To record overhead costs incurred)
9. i. Work in process inventory ($10600+3600)*200% $28400
Factory overhead $28400
(To record applied overhead to Jobs 137 and 140)

Calculation of Finished goods inventory

Finished goods inventory= Job 136+Job 138+Job 139

= $86600+132700+139600= $358900

Job 136= Direct materials+Direct labor+Factory overhead

Job 136= $50000+12200+(12200*200%)= $86600

Job 138= $19600+37700+(37700*200%)= $132700

Job 139= $22600+39000+(39000*200%)= $139600

Cost of goods sold= Job 136+Job 138

= $86600+132700= $219300

NOTE:- Please rate the answer and for any problem regarding the answer please ask in the comment section.